### v1

# Individualized and Interpretable Sleep Forecasting via a two-stage Adaptive Spatial-Temporal Model

Abstract
Sleep quality significantly impacts well-being. Therefore, healthcare providers and individuals need accessible and reliable forecasting tools for preventive interventions. This paper introduces an interpretable, individualized two-stage adaptive spatial-temporal model for predicting sleep quality scores. Our proposed framework combines multi-scale convolutional layers to model spatial interactions across multiple input variables, recurrent layers and attention mechanisms to capture long-term temporal dependencies, and a two-stage domain adaptation strategy to enhance generalization. The first adaptation stage is applied during training to mitigate overfitting on the training set. In the second stage, a source-free test-time adaptation mechanism is employed to adapt the model to new users without requiring labels. We conducted various experiments with five input window sizes (3, 5, 7, 9, and 11 days) and five prediction window sizes (1, 3, 5, 7, and 9 days). Our model consistently outperformed time series forecasting baseline approaches, including Long Short-Term Memory (LSTM), Informer, PatchTST, and TimesNet. The best performance was achieved with a three-day input window and a one-day prediction window, yielding a root mean square error (RMSE) of 0.216. Furthermore, the model demonstrated good predictive performance even for longer forecasting horizons (e.g, with a 0.257 RMSE for a three-day prediction window), highlighting its practical utility for real-world applications. We also conducted an explainability analysis to examine how different features influence sleep quality. These findings proved that the proposed framework offers a robust, adaptive, and explainable solution for personalized sleep forecasting using sparse data from commercial wearable devices.
I Introduction
Sleep plays a critical role in physical and mental health, and poor sleep quality is associated with a range of health risks, including cardiovascular disease, cognitive decline, and depression [1]. Most existing research on sleep assessment has focused on classification of sleep stages or retrospective sleep quality estimation, often relying on controlled laboratory data such as polysomnography (PSG) [2]. Studies with PSG provide high-fidelity insights, but they are costly, time-consuming, and impractical to scale for large populations. Furthermore, most studies do not address the predictive task of forecasting future sleep quality, which is essential for proactive health interventions. Moreover, in data sets that do only rely on a single night of recording, the first-night effect might alter symptoms.
Wearable devices represent a promising alternative to traditional polysomnography for sleep monitoring because they have expanded sleep forecasting applications beyond clinical environments, enabling personalized health monitoring. Wearable sensors such as accelerometers, photoplethysmography, and heart rate monitors embedded in commercial smartwatches like Garmin and Fitbit provide promising alternatives for unobtrusive long-term sleep monitoring, demonstrating significant reliability that makes them valuable tools for both personal health management and scientific research [3, 4]. Using data acquired by acti-watches, Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, a time-batched version of LSTM-RNN (TB-LSTM), these systems can analyze large volumes of data to achieve an operating characteristic (ROC) curve (AUC) of 0.9449 [5]. Moreover, wearable systems using objective sleep data have outperformed subjective self-reports in forecasting cognitive performance [6]. These innovations facilitate individualized feedback, early identification of sleep disorders [7], and population-level studies crucial for informing public health strategies. For instance, global smartphone-derived data has demonstrated how biological and cultural factors influence sleep patterns [8].
Several validation studies have compared the performance of smartwatches against gold-standard methods like polysomnography (PSG) and actigraphy. Fitbit devices, equipped with advanced sensors including accelerometers, optical heart rate monitors, and SpO2 sensors, have shown strong correlations with PSG in detecting sleep and wake periods, with studies reporting agreement rates over 85% for total sleep time and sleep efficiency [4, 9]. Commercial wearables like Fitbit and Garmin enable large-scale sleep research through affordability, cloud storage, and ease of use. Their non-invasive nature improves compliance over traditional PSG methods for longitudinal studies [10]. While not clinical-grade, these reliable devices advance sleep science and personalized health management. Deep learning models mentioned above [5] have achieved high predictive accuracy, but their inherent ”black-box” nature poses significant limitations in clinical utility and user trust. Especially in healthcare applications, interpretability is crucial for validating predictions, guiding treatment decisions, and ensuring fairness and reliability. Explainable AI (XAI) methodologies aim to bridge this gap by providing transparent insights into model decisions and elucidating the importance of various features [11].
Despite this abundance of wearable-generated data, there is a significant gap in the literature when it comes to forecasting sleep quality or sleep score trajectories, especially using data from commercial devices. In this work, we propose a novel adaptive, individualized, and explainable framework that bridges this gap by forecasting sleep quality scores using multimodal time-series data collected from a commercial smartwatch (Garmin Vivosmart 5). Our model is designed to generalize across users using a two-stage domain adaptation strategy and to operate effectively in real-world settings where labeled data are sparse and data distributions vary between individuals.
II Related Work
Contemporary research leverages multimodal, high-resolution data streams, including physiological signals such as EEG, heart rate variability, respiratory patterns, behavioral, and environmental data [12]. Despite such advances, individual differences and the complex interaction between sleep-wake cycles and sleep pressure create ongoing challenges [13].
Machine learning techniques have substantially improved the accuracy and utility of sleep forecasting. Deep learning models, such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and attention-based[14] architectures, excel at capturing temporal dependencies and modeling complex interactions in sleep data [15].
Nevertheless, generalization across diverse populations and environments remains challenging. Traditional machine learning models typically assume identical distributions between training and testing data, an assumption often violated in practical scenarios due to inter-subject variability and different data collection conditions. Transfer learning has enhanced the generalization capabilities of these models across diverse populations and heterogeneous data sources [16]. Moreover, domain adaptation techniques have emerged as a promising solution, facilitating learning of domain-invariant features and reducing distributional discrepancies [17]. Domain adaptation is crucial in sleep forecasting due to significant variations across populations, environments, and wearable technologies that cause models trained on specific groups or devices to generalize poorly. These variations include physiological differences in heart rate variability, circadian rhythms, and sleep architecture, environmental factors such as climate, noise, and lighting that influence sleep quality [18], and technological heterogeneity from different manufacturers’ measurement algorithms [6]. Domain adaptation methods address these challenges by transferring knowledge between domains (populations or devices), enabling personalized and reliable predictions essential for clinical and individual sleep management.
Figure 1: Illustration of domain adaptation for data from various subjects and conditions
TABLE I: Acronyms, units, ranges, anomaly, and missing data rates for variables from the Garmin Watch
Acronym Description Unit Anomaly Missing
TK TotalKilocalories kcal 15 (0.98%) 1 (0.09%)
TS TotalSteps arb. 56 (3.65%) 1 (0.09%)
TD TotalDistanceMeters meter 65 (4.23%) 1 (0.09%)
HA HighlyActiveSeconds seconds 34 (2.21%) 1 (0.09%)
AS ActiveSeconds seconds 27 (1.76%) 1 (0.09%)
MI ModerateIntensityMinutes minutes 51 (3.32%) 1 (0.09%)
RH RestingHeartRate beats per minute 59 (3.84%) 1 (0.09%)
MH minAvgHeartRate beats per minute 34 (2.21%) 2 (0.17%)
XH maxAvgHeartRate beats per minute 19 (1.24%) 2 (0.17%)
AWR AvgWakingRespirationValue breaths per minute 104 (6.67%) 2 (0.17%)
HRV HighestRespirationValue breaths per minute 51 (3.32%) 2 (0.17%)
LRV LowestRespirationValue breaths per minute 117 (7.62%) 2 (0.17%)
DS deepSleepSeconds seconds 25 (1.63%) 7 (0.61%)
LS lightSleepSeconds seconds 33 (2.15%) 7 (0.61%)
RS remSleepSeconds seconds 26 (1.69%) 40 (3.47%)
AW awakeSleepSeconds seconds 54 (3.52%) 7 (0.61%)
AC awakeCount arb. 27 (1.76%) 7 (0.61%)
$S$ avgSleepStress arb. 48 (3.12%) 7 (0.61%)
RM restlessMomentCount arb. 19 (1.24%) 2 (0.17%)
LR lowestRespiration breaths per minute 80 (5.21%) 6 (0.52%)
HR highestRespiration breaths per minute 65 (4.23%) 6 (0.52%)
AR averageRespiration breaths per minute 78 (5.08%) 6 (0.52%)
H Hydration milliliter N/A N/A
$S$ sleep_overallScore arb. 24 (1.56%) 7 (0.61%)
Test‐time adaptation (TTA) is an emerging inference‐time strategy that dynamically refines a pre‐trained model using incoming test data to address distributional shifts common in sleep forecasting. Unlike traditional models trained on fixed datasets—which often generalize poorly across age groups, health conditions, or different wearable devices—TTA enables on‐the‐fly adjustment of model parameters, improving personalization and accuracy without full retraining [19]. This is particularly valuable for handling inter‐ and intra‐individual variations in sleep patterns due to lifestyle changes or stress, as well as device heterogeneity arising from diverse data collection algorithms [20]. Moreover, by adapting in real time to environmental factors such as ambient noise or temperature fluctuations, TTA enhances the robustness and reliability of sleep‐quality predictions in real‐world settings [21].
Recent studies have integrated Explainable artificial intelligence (XAI) into sleep forecasting systems. In clinical settings, understanding contributing factors such as heart rate variability or environmental conditions enables clinicians to validate predictions against medical knowledge and make informed treatment decisions. For patients, interpretability increases engagement with behavioral interventions by clearly linking actions (e.g., screen time) to sleep outcomes [22, 23]. Additionally, XAI facilitates model improvement by identifying influential features and addressing biases, ensuring equitable and effective predictions across diverse populations [24]. Thus, XAI transforms sleep forecasting models from black-box predictors into transparent, actionable tools for personalized sleep management. Ahadian et al. [25] developed a trustworthy forecasting system employing Temporal Convolutional Networks (TCN), LSTM, and Temporal Fusion Transformers (TFT), complemented by SHAP-based counterfactual explanations and attention mechanisms. Their research shows that interpretability significantly improves model reliability, particularly for detecting sleep disorders across heterogeneous populations. Similarly, Göktepe-Kavis et al. [26] emphasize the necessity of explainable methods within AI-driven sleep scoring and diagnostic frameworks. They highlight the critical role interpretability plays in mitigating algorithmic biases and fostering clinical trust. Furthermore, Kumar et al. [27] stress the importance of interpretability for achieving standardized, generalizable AI models for sleep disorder detection. They argue that explainability is not only vital for user comprehension but also essential for regulatory compliance and ethical deployment in healthcare. Collectively, these studies highlight a growing consensus that explainability is indispensable in AI-driven sleep forecasting, particularly when transitioning predictive models from controlled laboratory settings to real-world clinical and personal healthcare applications.
III Dataset and pre-processing
III-A Dataset
This study was conducted in accordance with the Declaration of Helsinki and Dutch regulations governing research involving human participants. The Central Ethics Review Board for non-WMO studies (CTc UMCG), which oversees investigations outside the scope of the Medical Research Involving Human participants Act (WMO), reviewed and approved the study (study register number 18021). Written informed consent was obtained from each participant prior to enrollment.
Participants wore Garmin Vivosmart 5 devices for 10-15 weeks to record accelerations, heart rate, ambient light, and blood oxygen saturation, with instructions to remove the device daily as recommended by the manufacturer. They were recruited from two field labs in northern Netherlands: a weekly walking group supervised by healthcare professionals, and a 10-week personal lifestyle intervention program featuring personalized nutrition, cooking workshops, sports sessions, and mindfulness lessons. Research team members attended weekly walking sessions and visited the intervention site biweekly to provide device assistance, while participants otherwise maintained their regular routines. Of the 17 initial participants, one withdrew after three weeks, leaving 16 female participants with a median age of 55.5 years (IQR: 16.25) and median BMI of 28.22 kg/m² (IQR: 10.74), seven of whom had prior smartwatch experience. All data are maintained in the secure, offline Virtual Research Workspace (VRW) administered by the University of Groningen. Prior to downloading and local analysis, data are preprocessed within the VRW in an offline environment to remove sensitive information. More details of the dataset can be found in our previous work [28].
Figure 2: Boxplot for all features and sleep scores
III-B Anomaly and missing data
We implement a comprehensive multi-stage approach for handling anomalies and missing data in sleep quality prediction. The system first identifies missing data through two mechanisms: (i) invalid value detection, where sleep quality scores of -1 are replaced with NaN values, and (ii) zero value detection, where zero values in sleep quality scores are converted to NaN as they likely represent missing measurements. Missing and anomaly rates are summarized in Table I.
Figure 2 reveals numerous outliers across 24 Garmin smartwatch features that are physiologically implausible for participants aged over 50. Notable anomalies include totalSteps exceeding 30,000 (typical active range: <<10,000) and highlyActiveSeconds surpassing 20,000 (5.5 hours of high-intensity activity), both of which are highly unrealistic for this age demographic. The awakeSleepSeconds and restlessMomentCount features also show extreme values, suggesting sensor malfunction or improper device wear rather than genuine physiological measurements. These widespread anomalies across activity and sleep features necessitate robust anomaly detection before model training to prevent degraded sleep prediction performance and ensure reliable clinical insights. For anomaly detection, we employ two complementary approaches: a global statistical method using the interquartile range (IQR), which calculates Q1 (25th percentile) and Q3 (75th percentile) for each feature and defines anomalies as values outside [Q1 - 1×IQR, Q3 + 1×IQR], and a temporal deviation method that computes a 5-day rolling mean and identifies anomalies as values deviating more than 30 units from this mean. Following anomaly detection, we apply K-Nearest Neighbors (KNN) imputation with $k=3$ neighbors, which considers both cross-participant similarity and temporal patterns while preserving multivariate relationships between features. To address various types of noise in physiological time series data, we apply feature-specific smoothing methods based on physiological signal characteristics: exponential smoothing for heart rate (short-term trends), weighted moving average for activity (progressive patterns), adaptive smoothing for sleep stages (complex variability), Savitzky-Golay filtering for respiratory signals (feature preservation), and ensemble smoothing for the target sleep quality score (optimal robustness).
III-C Features
This study employs a single daily data point (i.e., sparse data) rather than original high-frequency measurements (one data point every 15 minutes) from Garmin, significantly reducing storage and computational demands. A total of 24 features, which were directly exported from the Garmin Connect App, were used in this study, including total kilocalories, total steps, total distance, highly active seconds, active seconds, moderate-intensity minutes, resting heart rate, minimum average heart rate, maximum average heart rate, average waking respiration value, highest respiration value, lowest respiration value, stress average, deep sleep seconds, light sleep seconds, REM sleep seconds, awake sleep seconds, awake count, average sleep stress, restless moment count, lowest respiration, highest respiration, average respiration, and sleep scores as shown in Table I and Figure 2. The predicting target, sleep score (0-100), is calculated using Firstbeat Analytic [29]. We removed the “Hydration” feature for all participants due to an excessive proportion of missing data because of manual input requirements, thereby ensuring a cleaner and more reliable data set. In addition, the first and last days of each participant’s recordings have been excluded due to insufficient device usage.
Selection Methods
We implement a comprehensive feature selection framework to identify the most relevant physiological and activity features for sleep quality prediction. Starting with 23 potential features encompassing activity metrics, heart rate, and respiration with cardio-respiratory metrics, respiration measurements, and sleep stage durations. We recognize that not all features contribute equally to prediction accuracy. Our framework allows dynamic feature selection through multiple methods, with the capability to reduce the feature set from 23 to 15 most informative features. We design this selection process to occur after data preprocessing but before model training, ensuring that only meaningful features are used for learning sleep quality patterns.
First, we computed Pearson correlation coefficients between each feature and sleep quality scores across all time points. We rank features by absolute correlation values and select those exceeding a threshold of 0.05, ensuring we capture both positive and negative relationships. If fewer than 15 features meet this threshold, we include additional high-correlation features to maintain our target dimensionality. Second, our Recursive Feature Elimination (RFE) approach uses a Random Forest regressor as the base estimator to iteratively remove the least important features. We configure RFE to eliminate one feature at a time until reaching our target of 15 features, providing rankings that reflect non-linear feature importance. By reducing from 24 to 15 features, we achieve approximately 32 percent dimensionality reduction, leading to faster training times, reduced overfitting risk, and improved model interpretability without sacrificing predictive performance.
We further implement mutual information-based selection to capture non-linear dependencies between features and sleep quality. This method computes the mutual information between each feature and the target variable, selecting the top 15 features with the highest information content. This approach proves particularly valuable for identifying features with complex, non-monotonic relationships to sleep quality. Our most sophisticated approach is the ensemble feature selection method, which combines insights from all three previous methods. We run correlation, RFE, and mutual information selection with relaxed constraints (selecting top 15 features from each), then implement a voting mechanism where features receive votes based on their selection by each method. We select the 15 features with the highest vote counts, ensuring robust selection that is not biased by any single method’s assumptions.
We validate our feature selection approach through multiple mechanisms. We compare model performance with and without feature selection, consistently observing comparable or improved prediction accuracy with reduced feature sets. Our interpretability analysis using SHAP values to confirm that selected features align with domain knowledge about sleep physiology. The consistency of selected features across different cross-validation folds demonstrates the stability of our selection process.
III-D Normalization
We applied a MinMaxScaler from sklearn, scaling the feature data to a standard range (0 to 1) for train, validation, and test sets separately to avoid data leak. Each feature is normalized before being fed into the model. The predicted results are then denormalized to compare them with the true values and calculate the loss.
III-E Sliding window
To pre-process the time series data, we employ a sliding window approach with window size $w$ (3, 5, 7, 9, 11 days), stride $s$ (one day), and prediction window size $\delta$ (1, 3, 5, 7, 9 days), transforming each participant’s dataset into a set of windowed instances for data augmentation as shown.
Figure 3: Domain shifts for features from different subjects by PCA.
III-F Evaluation metrics
We use Mean Square Error (MSE), Mean Absolute Error (MAE), and Root Mean Square Error (RMSE) as metrics in this work because they are frequently adopted in forecasting tasks, particularly in deep learning contexts, due to their interpretability and computational simplicity. They are defined as:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|.$
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
where $y_{i}$ denotes the normalized sleep value on day $i$, $\hat{y}_{i}$ represents the corresponding predicted normalized value, and $n$ is the total number of days.
III-G Leave-one-out cross-validation
In many machine learning problems involving human participants, data collected from different individuals (participants) may vary significantly due to physiological, behavioral, or demographic factors. One standard approach for evaluating the robustness of a model across different individuals is leave-one-out cross-validation (LOOCV). This approach ensures that for each participant, the model is evaluated on data that remain completely unseen during training, providing a realistic estimate of cross-participant generalization performance. There are 16 distinct participants in this work, each associated with a dataset:
$\mathcal{S}_{p}=\{(\mathbf{f}_{p,k},q_{p,k})\mid k=1,2,\ldots,T_{p}\},$ (1)
where:
  • $\mathbf{f}_{p,k}\in\mathcal{F}$ is the feature vector of the $k$-th sample from participant $p$,
  • $q_{p,k}\in\mathcal{Q}$ is the sleep quality score corresponding to $\mathbf{f}_{p,k}$,
and $T_{p}$ is the total number of samples from participant $p$. Hence, the entire dataset by 16 participants in this work can be written as
$\mathcal{D}=\bigcup_{s=1}^{16}\mathcal{D}_{s}.$ (2)
In this work, we split the data by participant. Concretely, for each participant $s\in\{1,2,\ldots,16\}$: Practitioners split the training data (from the $N$ subjects) into a separate training set (N2N-2 participants), a validation set (one participant) and a test set (one participant). For instance,
$\text{Train set}=\mathcal{D}_{a},\quad\text{Validation set}=\mathcal{D}_{b},\quad\text{Test set}=\mathcal{D}_{s},$ (3)
where $a\neq s$, $b\neq s$, and $a\neq b$. Hyperparameters are tuned on $\mathcal{D}_{b}$ (validation set), while ultimately evaluating performance on the unseen participant $\mathcal{D}_{s}$ (test set).
TABLE II: Performance comparison of RMSE for various models and input and prediction window sizes
Model Input = 3 days Input = 5 days Input = 7 days
1 3 5 7 9 1 3 5 7 9 1 3 5 7 9
LSTM [30] 0.300 0.304 0.298 0.316 0.299 0.312 0.299 0.305 0.297 0.310 0.332 0.308 0.336 0.335 0.321
Informer [31] 0.290 0.290 0.290 0.293 0.290 0.293 0.297 0.294 0.292 0.298 0.296 0.294 0.294 0.293 0.293
PatchTST [32] 0.292 0.294 0.296 0.297 0.299 0.285 0.294 0.301 0.299 0.299 0.299 0.291 0.308 0.303 0.317
TimesNet [33] 0.283 0.294 0.297 0.329 0.322 0.279 0.293 0.305 0.303 0.330 0.294 0.292 0.294 0.307 0.312
Ours 0.216 0.257 0.272 0.316 0.317 0.231 0.234 0.229 0.317 0.318 0.279 0.286 0.249 0.313 0.320
Model Input = 9 days Input = 11 days
1 3 5 7 9 1 3 5 7 9
LSTM [30] 0.300 0.305 0.333 0.320 0.344 0.289 0.308 0.304 0.315 0.312
Informer [31] 0.292 0.292 0.295 0.294 0.291 0.290 0.292 0.294 0.292 0.291
PatchTST [32] 0.279 0.302 0.295 0.321 0.328 0.307 0.310 0.317 0.310 0.308
TimesNet [33] 0.291 0.299 0.313 0.319 0.314 0.285 0.303 0.309 0.315 0.322
Ours 0.307 0.311 0.267 0.306 0.317 0.265 0.287 0.297 0.321 0.313
After performing this procedure for each participant $i$, we obtain a set of metrics on the $i$-th fold: MAE $\{MAE_{1},\ldots,MAE_{N}\}$, MSE $\{MSE_{1},\ldots,MSE_{N}\}$, and RMSE $\{RMSE_{1},\ldots,RMSE_{N}\}$. Leave-one-out cross-validation (LOOCV) is a rigorous strategy for evaluating participant-based machine learning tasks. By training on N2N-2 participants , validating on another participant and testing on the remaining participant, we obtain an unbiased estimate of cross-participant generalization.
Algorithm 1 Our model with Two-Stage Domain Adaptation
Input: $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{test}}$, Epochs $E$, Domain weight $\alpha$ Output: Trained model $\mathcal{M}$
1:  Phase 1: Training with Domain Adaptation
2:  Initialize $\mathcal{M}$ (Multi-scale CNN + BiLSTM + Self-Attention + Domain Classifier)
3:  Initialize Adam optimizer with $\text{lr}=10^{-3}$
4:  for epoch =1=1 to $E$ do
5:   for batch $(X,y,d)\in\mathcal{D}_{\text{train}}$ do
6:    $\hat{y},\hat{d}\leftarrow\mathcal{M}(X,\text{return\_domain}=\text{True})$
7:    $\mathcal{L}\leftarrow\text{MSE}(\hat{y},y)+\alpha\cdot\text{CrossEntropy}(\hat{d},d)$
8:    Update $\mathcal{M}$ via backpropagation
9:   end for
10:   Evaluate on $\mathcal{D}_{\text{val}}$ and apply Early Stopping
11:  end for
12:  Phase 2: Test-Time Adaptation
13:  Initialize TTA optimizer with $\text{lr}_{\text{tta}}=10^{-4}$
14:  for epoch =1=1 to $E_{\text{tta}}$ do
15:   for batch $(X_{\text{test}},-,-)\in\mathcal{D}_{\text{test}}$ do
16:    Generate $X_{\text{aug1}},X_{\text{aug2}}$ with noise augmentation
17:    $\hat{y}_{\text{orig}}\leftarrow\mathcal{M}(X_{\text{test}})$, $\hat{y}_{\text{aug1}}\leftarrow\mathcal{M}(X_{\text{aug1}})$, $\hat{y}_{\text{aug2}}\leftarrow\mathcal{M}(X_{\text{aug2}})$
18:    $\mathcal{L}_{\text{tta}}\leftarrow\frac{1}{3}\sum\text{MSE}(\hat{y}_{i},\hat{y}_{j})$ for all pairs
19:    Update $\mathcal{M}$ via backpropagation
20:   end for
21:  end for
22:  Phase 3: Final Inference
23:  $\mathcal{M}.\text{eval}()$
24:  for batch $(X_{\text{test}},y_{\text{test}},-)\in\mathcal{D}_{\text{test}}$ do
25:   $\hat{y}_{\text{final}}\leftarrow\mathcal{M}(X_{\text{test}})$
26:  end for
27:  Compute RMSE, MAE, MSE
IV Methodology
IV-A Models
We propose a spatial-temporal model with attention mechanisms as our primary architecture for sleep quality prediction, as shown in the flowchart of Figure 4. This model was specifically designed to address the unique challenges of multi-modal physiological time series data, which exhibits both local patterns (e.g., short-term heart rate variations) and long-range temporal dependencies (e.g., sleep stage transitions across nights). Our architecture integrates multiple state-of-the-art components: a multi-scale CNN module with parallel convolutional layers using kernel sizes of {3, 5, 7} plus dilated convolutions, a channel attention mechanism for automatic feature selection, bidirectional LSTM layers for capturing temporal dependencies in both directions, and dual attention mechanisms including 8-head self-attention and temporal attention for identifying crucial time points. These components work synergistically to extract meaningful patterns from complex physiological signals.
We chose this hybrid architecture based on the complementary strengths of convolutional and recurrent networks for physiological time series analysis. Multi-scale CNNs excel at extracting local temporal patterns and are computationally efficient through weight sharing, while BiLSTMs capture long-term dependencies and handle variable-length sequences naturally. The attention mechanisms address the limitation of traditional RNNs in accessing distant temporal information, while the multi-scale approach acknowledges that sleep-related features manifest at different time scales [34].
IV-B Domain Adaptation
In practical scenarios, the data distribution of the training set (source domain) may differ from that of the test set (target domain) due to inter-subject variability. One can clearly find domain shifts of data from each participant due to individual aspects like gender, age, lifestyle, etc, as shown in Figure 3. Figure 3 reveals significant domain shift in our dataset through PCA visualization of 16 subjects. Each participant’s data forms distinct, minimally-overlapping clusters in the reduced feature space, with subjects distributed across different regions—for example, Subject 3 occupies the lower-left quadrant while Subject 4 clusters in the upper-right. The convex hulls emphasize this separation, with some subjects (S1, S8, S10) showing compact clusters and others (S3, S7, S14) exhibiting dispersed patterns. This substantial inter-subject variability in physiological and behavioral patterns demonstrates that traditional assumptions fail for this data, necessitating domain adaptation techniques to develop robust sleep prediction models that generalize across diverse individual profiles.
This distributional shift can degrade the model’s performance on unseen participants. $\mathcal{D}_{a}=\mathcal{D}_{\text{train}}$ represent the source domain and $\mathcal{D}_{t}=\mathcal{D}_{q}$ denote the target domain for a given fold $(i,q)$. The joint distributions of the source and target domains are:
$P_{s}(\mathbf{x},\mathbf{y})\;\neq\;P_{t}(\mathbf{x},\mathbf{y}),$
where $P_{s}$ and $P_{t}$ are the probability distributions over $\mathcal{X}\times\mathcal{Y}$ for the source and target domains, respectively.
Domain adaptation seeks to mitigate this issue by aligning the source and target domains, thereby enhancing the model’s generalization ability. When distributions vary significantly across participants (due to physiology, behavior, etc.), each participant can be viewed as its own domain. Training on $\mathcal{D}_{p}$, validating on $\mathcal{D}_{q}$, and testing on $\mathcal{D}_{i}$ can be interpreted as adapting from a source domain to a target domain. This viewpoint motivates the application of domain adaptation methods, which aim to reduce distributional discrepancies between training participants and a new target participant.
Our Domain Adaptation Strategy
Recognizing that physiological patterns vary significantly between individuals, we designed our framework to ensure that models trained on one group of participants can effectively generalize to new, unseen participants. We implement a two-stage domain adaptation approach to address the critical challenge of cross-participant generalization in sleep quality prediction as shown in the pseudo code from Algorithm 1. Our adaptation strategy operates in two distinct stages: training-time domain adaptation and test-time adaptation (TTA). We provide flexible configuration options allowing these stages to be used independently or in combination through four modes: ’none’, ’train-only’, ’test-only’, or ’both’, enabling systematic evaluation of each component’s contribution.
Stage 1: Training-Time Domain Adaptation
In the first stage, we implement domain adaptation during model training through adversarial learning. We equip each model architecture with a domain classifier component designed to distinguish between different participants (domains). Our training process employs a minimax game where the feature extractor learns representations that are simultaneously predictive of sleep quality and invariant across participants. We formulate the total loss as a combination of the primary sleep quality prediction loss and a domain classification loss, controlled by a weighting parameter alpha. Through gradient reversal, we train the domain classifier to identify the source participant while simultaneously training the feature extractor to produce representations that confuse the domain classifier. This adversarial training mechanism encourages our models to learn participant-invariant features that enhance generalization to new individuals.
Stage 2: Test-Time Adaptation (TTA)
We introduce the second adaptation stage during inference when our model encounters data from previously unseen participants. We implement three complementary TTA strategies. First, our entropy-based TTA leverages the model’s prediction confidence for self-supervised adaptation. We identify high-confidence predictions (using a threshold of 0.9) and employ them as pseudo-labels to fine-tune the model on the test participant’s unlabeled data. Second, our consistency-based TTA enforces prediction stability under input perturbations. We update the model to minimize discrepancies between predictions on original inputs and their augmented versions (with controlled noise levels of 0.01 and 0.02). Third, we develop temporal consistency TTA that exploits the sequential nature of sleep data by enforcing smooth predictions across time and maintaining consistency between temporally shifted sequences.
Implementation Design and Flexibility
We designed our framework with flexibility as a core principle, allowing systematic investigation of each adaptation component’s effectiveness. We enable four operational modes: setting mode to ’none’ provides baseline performance without adaptation, ’train-only’ isolates the effect of adversarial training during the learning phase, ’test-only’ evaluates the impact of test-time adaptation alone, and ’both’ combines both stages for comprehensive adaptation. We configure training-time adaptation with a domain loss weight of 1.0 and implement gradient reversal for effective adversarial feature learning. For test-time adaptation, we employ 20 adaptation epochs with a learning rate of $1\times 10^{-4}$, with the specific strategy determined by the tta-method parameter.
Rationale and Empirical Benefits
The training-time adaptation ensures our models learn robust, generalizable features across the training population, mitigating overfitting to individual participants’ idiosyncratic patterns. The test-time adaptation enables personalization when deploying the model to new participants, allowing dynamic adjustment to individual physiological characteristics without requiring additional labeled data. Our experimental results validate this design through systematic comparison across adaptation modes. The pseudo code is shown above in the info-graphic Algorithm 1.
IV-C Baseline Models for Comparative Analysis
To evaluate our proposed architecture, we implemented several baseline models representing different modeling paradigms. The LSTM Model uses stacked LSTM layers without convolutional preprocessing, evaluating the importance of explicit local feature extraction. We also compared against recent transformer-based architectures that have shown promise in time series forecasting, including PatchTST [32], Informer [31], and timesnet [33].
IV-D Explainable AI for sleep forecast
Explainable Artificial Intelligence (XAI) enhances trust and adoption by making model predictions transparent and interpretable for healthcare professionals and patients alike [35]. We implement SHAP analysis [36] using the model-agnostic KernelExplainer to interpret our deep learning models. To handle the three-dimensional input data (samples × time steps × features), we apply temporal aggregation, using 50 background samples and 100 test samples for computational efficiency. We calculate SHAP values through the standard Shapley value formula [36]. This approach provides both feature rankings and directional insights, enabling clinicians to understand not only which factors matter but how they influence sleep quality. The framework supports personalized explanations through individual SHAP values while maintaining computational tractability through temporal aggregation and strategic sampling.
IV-E Training Procedure
Our training methods consist of two optional phases. In Phase 1, the model could be trained with domain adaptation using a combined loss function $\mathcal{L}=\mathcal{L}_{\text{main}}+\alpha\cdot\mathcal{L}_{\text{domain}}$, where $\mathcal{L}_{\text{main}}=\text{RMSE}(\hat{y},y)$ represents the sleep quality prediction loss and $\mathcal{L}_{\text{domain}}=\text{CrossEntropy}(\hat{d},d)$ enforces domain-invariant feature learning across participants, with the domain weight $\alpha$ optimized during hyperparameter tuning. The training employs Adam optimizer with learning rate $10^{-3}$, which incorporates early stopping with smoothed validation loss (patience=30, min_delta=0.0001), and applies comprehensive data preprocessing including ensemble smoothing methods and automated feature selection to 15 most relevant features from the original 23 Garmin metrics. In Phase 2: we implement Test-Time Adaptation (TTA) on the held-out test participant using consistency regularization with augmented inputs, minimizing $\mathcal{L}_{\text{TTA}}=\frac{1}{3}\sum_{i,j}\text{RMSE}(\hat{y}_{i},\hat{y}_{j})$ across original and noise-augmented predictions with learning rate $10^{-4}$ for 10 epochs. Phase 3 conducts final inference on the adapted model, and this entire procedure is repeated for each participant in the LOOCV framework, ensuring robust evaluation across individual differences while the domain adaptation components help the model generalize across different participants’ physiological patterns and behavioral variations in sleep data.
Figure 4: Flowchart of our adaptive spatial-temporal model with optional two-stage domain adaptation
IV-F Hyperparameter Optimization
All models in this work were optimized using Optuna’s Tree-structured Parzen Estimator (TPE) sampling within a Leave-One-Out cross-validation (LOOCV) framework to ensure robust generalization across participants. The search space included architectural parameters: number of convolutional layers $\in\{1,2\}$, LSTM layers $\in\{1,2,3\}$, CNN hidden dimensions $\in\{16,32,64\}$, LSTM hidden dimensions $\in\{64,128,256\}$, and regularization parameters: CNN dropout rate $\in[0.1,0.5]$, LSTM dropout rate $\in[0.1,0.5]$, batch normalization toggle $\in\{\text{True},\text{False}\}$, batch size $\in\{8,16,32\}$, and domain adaptation weight $\alpha\in[0.0,1.0]$. The optimization objective minimized the average RMSE across all LOOCV folds: $\theta^{*}=\arg\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\text{RMSE}(\mathcal{M}_{\theta},S_{i})$, where $\mathcal{M}_{\theta}$ represents the our model parameterized by $\theta$, and $S_{i}$ denotes the $i$-th held-out participant. Early stopping with smoothed validation loss (smoothing factor $\beta=0.1$) was applied with patience of 30 epochs and minimum improvement threshold of 0.0001 to prevent overfitting during the hyperparameter search process.
IV-G Hardware and Software Setup
Experiments were conducted on two hardware platforms: a local workstation (Intel i7-13700K, NVIDIA RTX 4090, 128 GB RAM) for test and development, and the Hábrók HPC cluster [^1] https://wiki.hpc.rug.nl/habrok/introduction/cluster_description for large-scale computations like hyperparameter fine-tuning. The cluster comprises 190 heterogeneous nodes including standard compute nodes (119×128 AMD EPYC cores), high-memory nodes (4×80 Intel Xeon cores, 4 TB RAM), and GPU nodes featuring NVIDIA A100, V100, L40s, and H100 accelerators. All models were implemented in PyTorch for cross-platform consistency.
V Results and Discussion
Our approach captures both local feature interactions and longer-term temporal dependencies, while a two-stage adaptation provides adaptation across different participants. Our model offers several critical advantages for sleep quality prediction. First, the multi-scale feature extraction capability allows the model to simultaneously capture patterns at different temporal resolutions—from minute-level heart rate fluctuations to hour-level sleep stage progressions. Second, the attention mechanisms provide interpretability by highlighting which features and time points most influence predictions, crucial for clinical applications. Third, the bidirectional processing leverages both historical context and future patterns within the input window, particularly valuable for understanding sleep preparation and recovery phases. Fourth, residual connections between CNN and output layers ensure stable gradient flow during training while preserving fine-grained temporal information. Finally, the integrated domain adaptation components enable robust cross-participant generalization, addressing the significant inter-individual variability in sleep patterns. We observe consistent improvements across all metrics (MAE, MSE, and RMSE) when employing domain adaptation. The most substantial gains are achieved when both adaptation stages are active (mode=’both’), yielding an RMSE of 0.216 compared to 0.245 for training-only adaptation and 0.242 for test-only adaptation. This finding confirms our hypothesis that combining population-level robustness (from training-time adaptation) with individual-level personalization (from test-time adaptation) provides optimal performance for cross-participant sleep quality prediction. Our approach is particularly valuable in real-world healthcare applications where physiological variability is substantial and obtaining labeled data from every new patient is often impractical.
V-A Evaluation with Different Window Sizes
We tested five input window sizes (3, 5, 7, 9, and 11 days) in combination with five predicting window sizes (1, 3, 5, 7, 9 days) on both baseline models and our proposed adaptive spatial and temporal model. Table II summarizes the results. From the results, it can be observed that all models, including our proposed approach, generally achieve their lowest RMSE when the prediction window size is one day. Longer prediction windows are also possible with the increase in errors. Our model achieves the best performance compared to other base models and state-of-the-art time series forecasting models.
Figure 5: True values vs predictions in validation and test set for subjects.
V-B Temporal Prediction Analysis and Trend Accuracy
Figure 5 presents a comparison between ground-truth sleep scores and model predictions across both validation and test sets. To maintain clarity and avoid redundancy, we display representative participants that demonstrate the range of model performance patterns observed across various participants. The visualizations reveal that our model not only accurately predicts overall sleep quality levels but also effectively captures day-to-day variations and long-term sleep pattern changes for most participants.
To comprehensively evaluate model performance, we simultaneously display raw predictions and smoothed trend lines. Raw data is shown with faded colors, while trend lines are highlighted using darker colors and thicker lines. Validation and test sets are distinguished through different color schemes: blue-orange for validation set and blue-red for test set. By calculating rolling standard deviations, we add confidence bands to each prediction using semi-transparent fills to represent prediction uncertainty ranges. This visualization method intuitively demonstrates the stability of model predictions. We separately calculate Pearson correlation coefficients for validation and test sets to evaluate the model’s ability to capture sleep quality trends: Beyond traditional regression metrics, we introduce trend direction accuracy to evaluate the model’s ability to predict sleep quality change directions.
Through trend correlation assessment, we can verify whether the model captures long-term sleep pattern changes, which is crucial for monitoring chronic sleep disorders. Trend direction accuracy directly reflects the model’s ability to predict sleep quality improvement or deterioration, providing important guidance for timely interventions. By displaying prediction range bands, clinicians can understand prediction confidence levels, facilitating more cautious decision-making. Our visualization method uses color coding (green for correct direction predictions, blue for incorrect) to display model performance at each time point, enabling quick identification of model strengths and weaknesses.
Using rolling window methods to calculate trends and standard deviations improves evaluation robustness and reduces outlier effects. The visualization in Figure 5 reveals a strong alignment between predicted and actual sleep scores across most participants. Several key observations emerge:
Trend Correlation Performance: The trend correlation values demonstrate robust model performance, test set correlations spanning 0.794 (participant 6) to 0.932 (participant 9). The consistently high correlation values (most exceeding 0.88) indicate that our model effectively captures the underlying sleep quality patterns across diverse individuals.
Trend Direction Accuracy: The trend direction accuracy metric shows performance, with test set accuracies ranging from 67.1% (participant 1) to 86.2% (participant 9). These high accuracy rates demonstrate the model’s ability to correctly predict whether sleep quality will improve or deteriorate, which is crucial for clinical intervention timing.
V-B1 Individual participant Analysis
The participant-specific patterns reveal interesting variations in model performance. Participants (1, 4, 7, 9, 11) exhibit exceptional model fit with both high trend correlations (>>0.85) and direction accuracies (>>80%). The visual alignment between predicted and actual values is particularly strong, with the model successfully capturing both short-term fluctuations and long-term trends. While still achieving respectable performance metrics, participants (4, 7, 11) show slightly more prediction variance. Participant 7, for instance, demonstrates good test set performance (90.3% direction accuracy) despite having the lowest validation correlation (0.561), suggesting successful model adaptation during testing. Participants with highly variable sleep patterns (e.g., participant 11’s dramatic fluctuations around day 100) present greater prediction challenges, yet the model maintains reasonable accuracy by correctly capturing the general trend direction even when exact values differ.
V-B2 Temporal Dynamics
The temporal analysis reveals several important characteristics:
Validation vs. Test Set Transition: The gray vertical line separating validation and test sets shows generally smooth transitions in prediction quality, indicating robust model generalization. The absence of significant performance degradation in test sets validates our training methodology.
Uncertainty Quantification: The shaded regions representing prediction ranges effectively bound the actual values in most cases, with wider uncertainty bands appearing during periods of higher sleep variability. This adaptive uncertainty quantification provides valuable confidence information for clinical decision-making.
Direction Prediction Visualization: The color-coded bars at the bottom of each plot (green for correct direction, blue for incorrect) provide immediate visual feedback on prediction reliability. Notably, clusters of correct predictions often coincide with stable sleep patterns, while incorrect predictions tend to occur during rapid transitions.
V-B3 Clinical Relevance
The high trend direction accuracies across participants have clinical implications:
  • Early Intervention: With direction accuracies consistently above 75%, clinicians can confidently use model predictions to identify when a patient’s sleep quality is likely to deteriorate, enabling proactive interventions.
  • Personalized Prediction: The variation in model performance across participants (e.g., participant 9’s 0.932 test correlation vs. participant 6’s 0.794) suggests that some individuals have more predictable sleep patterns, allowing for customized monitoring frequencies.
  • Trend-Based Prediction: The strong trend correlations indicate that the model captures meaningful sleep quality evolution rather than just noise.
V-B4 Model Robustness
The analysis demonstrates several aspects of model robustness:
  • Consistency Across participants: Despite individual sleep pattern variations, the model maintains high performance metrics across all participants.
  • Temporal Stability: The model performs consistently across extended time periods (up to 170 days for some participants), suggesting good long-term reliability.
  • Adaptation to Patterns: The model successfully adapts to different sleep pattern types, from highly regular (participant 6) to more variable (participant 11), demonstrating flexible feature learning.
These results collectively validate our model’s clinical utility for continuous sleep quality prediction, providing both accurate point estimates and reliable trend information essential for effective sleep disorder management.
Figure 6: Comparison of ALL participants’ MAE, MSE, and RMSE by different models in terms of three-day input and one-day prediction window size.
V-C Cross-participant Performance Comparison
Figure 6 presents a comprehensive comparison of model performance across all participants using three key metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE) in terms of three-day input and one-day prediction window size. The radar plots enable simultaneous visualization of individual participant performance and overall model behavior patterns.
The shorter radial distances for our model indicate substantially lower error compared to the baselines in almost every case. This consistency suggests that our spatial and temporal model with two-stage domain adaptation techniques enables our approach to robustly handle participant-specific variability and produce accurate sleep forecasts.
Overall, these findings validate the effectiveness of the proposed method. The model not only excels with shorter predicting windows (one day) but also sustains competitive accuracy for extended horizons. Such versatility could be valuable for real-world applications where sleep monitoring and prediction over multiple days can inform better health interventions and personalized recommendations.
V-C1 Model Performance Patterns
The radar plots reveal several important patterns in model performance:
Overall Performance Hierarchy: Across all three metrics, a consistent performance ranking emerges, with our proposed model demonstrating superior performance (smaller error areas) compared to baseline methods. Notable performance variations exist across participants: participants 1, 5, 6, and 13 show relatively low error values across all models, suggesting these participants have more predictable sleep patterns. It exhibits higher error values, particularly pronounced in the MSE and RMSE metrics, indicating greater prediction challenges for participants 0, 10, and 12. Participant 7 demonstrates exceptional model performance with minimal errors across all metrics, suggesting highly regular sleep patterns.
V-C2 Metric-Specific Observations
The MSE plot (left panel) shows the error values ranging from near 0 to 0.25. The quadratic nature of MSE amplifies the performance gaps, making it particularly useful for identifying participants where models struggle. Our model maintains the smallest area coverage, indicating consistent superiority in minimizing large prediction errors. The MAE plot (center panel) provides a more balanced view of average prediction accuracy, with values ranging from 0 to 0.4. The linear nature of MAE shows that while absolute differences between models are smaller than in MSE, the relative performance ranking remains consistent. This suggests that our model’s advantages are not merely due to outlier handling but represent genuine improvements in average prediction accuracy. The RMSE plot (right panel) combines the sensitivity of MSE to large errors with the interpretability of MAE’s scale. Values range from 0 to 0.5, and the patterns closely mirror those in the MSE plot but with a compressed scale. This metric confirms that our model’s superior performance is robust across different error measurement approaches.
Our Model (Red) consistently maintains the smallest error region across all metrics, with particularly strong performance on challenging participants (0, 8, 9, 10, 11, 12, 13, 14, 15). The compact radar area indicates robust generalization across diverse sleep patterns.
These results validate our model’s effectiveness in handling the inherent variability in sleep patterns across individuals while maintaining superior prediction accuracy compared to state-of-the-art baselines.
Figure 7: Features importance analysis with impact direction by SHAP analysis
V-D Explainable model for sleep forecast
The results reveal distinct patterns of feature importance across participants, demonstrating both individual variability and common influential factors in sleep quality prediction.
Average Feature Importance Analysis: The aggregated analysis of all 16 subjects reveals that deep sleep seconds (DS) emerges as the strongest positive predictor of sleep quality. This finding validates the fundamental importance of deep sleep stages in determining overall sleep quality and restorative function. The robust positive impact of DS across the entire subject population underscores the critical role of slow-wave sleep in sleep quality assessment. Additionally, resting heart rate (RH) demonstrates a consistent moderate positive influence (mean $\approx$ +0.4), reinforcing the relationship between cardiovascular fitness and sleep quality outcomes.
V-D1 Clinical Implications of Feature Importance
The SHAP analysis provides several clinically relevant insights:
  1. 1.
    Restlessness as a Primary Indicator:The moderate positive influence of sleep stage durations (Deep Sleep Seconds, Light Sleep Seconds, Rem Sleep Seconds) confirms that our model appropriately values sleep architecture. The balanced contributions suggest the model considers overall sleep composition rather than overemphasizing any single stage.
  2. 2.
    Individual Variability: The individual subject analyses reveal substantial heterogeneity in feature response patterns, with response scores ranging from Subject 13’s exceptional restless moment count (RM) impact of +5.0+5.0 to Subject 03’s severe awake count (AC) sensitivity of 2.5-2.5. Subject 03 exhibits the most extreme sensitivity to sleep fragmentation, with AC demonstrating an exceptionally strong negative impact (2.5-2.5) while RM shows a strong positive impact (+1.8+1.8), and resting heart rate (RH) unexpectedly shows a negative impact (0.8-0.8), contrasting with the general population pattern. Subject 06 demonstrates a unique cardiovascular-sleep interaction pattern, with RH showing the strongest negative impact (1.0-1.0) among all measured features, while RM exhibits a high positive impact (+1.4+1.4), suggesting interventions targeting both cardiovascular health and sleep environment optimization would be beneficial. Subject 13 presents the most remarkable profile, with RM demonstrating an unprecedented positive impact approaching +5.0+5.0, indicating that natural sleep movements are critically important for this individual’s sleep quality maintenance. Subject 14 exhibits a balanced response pattern with AC showing substantial negative impact (1.4-1.4), while RM and deep sleep seconds (DS) both demonstrate strong positive influences (+1.0+1.0 each), and total steps (TS) shows notable negative impact (0.7-0.7). Subject 16 demonstrates a distinctive activity-responsive pattern where total kilocalories (TK), highly active seconds (HA), and RM all exhibit substantial positive impacts (+0.8+0.8, +0.8+0.8, and +0.7+0.7 respectively), while AC shows strong negative influence (1.0-1.0) and both RH and TS show moderate negative impacts.
    The differences in feature response patterns across subjects underscore the critical importance of personalized sleep medicine approaches, as this 7.5-point range in individual feature responses far exceeds population-level variations observed in the aggregated analysis. The identification of distinct phenotypes—including sleep fragmentation sensitive (Subject 03), cardiovascular-arousal sensitive (Subject 06), movement-dependent (Subject 13), multi-factor balanced (Subject 14), and activity-responsive (Subject 16) profiles—provides a framework for developing targeted interventions.
V-D2 Model Behavior Validation
The interpretability results demonstrate several positive aspects of our model’s behavior:
  1. 1.
    Physiological Coherence: The model assigns importance to physiologically meaningful features. Respiratory values during sleep and waking states show consistent but moderate importance, indicating the model captures subtle physiological patterns without overfitting to noise.
  2. 2.
    Absence of Spurious Correlations: Features with low clinical relevance are appropriately de-emphasized, suggesting the model has learned meaningful rather than spurious relationships.
  3. 3.
    Bidirectional Effects: The presence of both positive and negative SHAP values for different features indicates the model captures complex, relationships. For instance, while increased deep sleep positively influences predictions, excessive restlessness negatively impacts them, reflecting real-world sleep dynamics.
V-D3 Practical Applications
The interpretability framework enables several practical applications:
  1. 1.
    Personalized Sleep Interventions: By identifying individual-specific important features, clinicians can tailor interventions. The high ranking of daily activity in participant 03, for example, could indicate that an intervention targeting activity might be more beneficial for this individual.
  2. 2.
    Model Trust and Adoption: The ability to explain predictions using clinically meaningful features enhances trust in the model’s recommendations, facilitating adoption in clinical settings.
  3. 3.
    Feature Engineering Guidance: The importance rankings can guide future feature engineering efforts, potentially focusing on more detailed restlessness metrics or sleep fragmentation indices.
This interpretability analysis transforms our deep learning model from a “black box” into a transparent tool that provides actionable insights for both clinicians and patients.
VI Conclusions and Limitations
This study presents a comprehensive deep learning framework for sleep quality forecasting that addresses key challenges in explainable and personalized sleep management. Our two-stage domain adaptation model consistently outperforms state-of-the-art baselines across all evaluation metrics, achieving trend correlations exceeding 0.8 for most participants and trend direction accuracies above 75%. These results demonstrate the model’s robust capability in capturing both short-term fluctuations and long-term sleep quality patterns.
Our model could inform users and clinicians about future potential deterioration, thereby facilitating individualized and timely interventions. The model’s ability to maintain performance across extended periods (up to 170 days) and diverse sleep patterns validates its potential for real-world deployment in continuous sleep monitoring systems. Through SHAP analysis, we identified Deep Sleep Seconds as the most influential predictor across participants, while revealing significant individual variations in feature importance. This interpretability framework transforms our model from a black box into a transparent tool that provides actionable insights for personalized sleep interventions.
Our model demonstrates remarkable consistency across 16 participants with diverse sleep patterns, from highly regular to extremely variable. The smooth transition between validation and test sets, coupled with adaptive uncertainty quantification, confirms the model’s generalization capability and clinical reliability. Furthermore, our findings emphasize that effective sleep quality prediction requires individualized approaches, as substantial inter-participant variations in both model performance and feature importance highlight the inadequacy of one-size-fits-all solutions in sleep management.
Despite promising results, several limitations merit consideration. First, the model relies solely on wearable sensor data without incorporating environmental factors (temperature, noise), lifestyle variables (caffeine, work schedules), or psychological states that could enhance prediction accuracy for irregular patterns. Second, while demonstrating strong technical performance, prospective clinical validation is needed to assess real-world impact on sleep interventions. The study is limited by using Garmin-derived sleep scores that may not represent a clinically recognized sleep quality measure, though the approach is applicable to other sleep assessment methods. Future work should focus on adaptive models for sudden pattern changes, continuous learning frameworks, multi-modal data fusion, and standardized benchmarks to advance toward clinically impactful solutions.
ACKNOWLEDGMENT
The work is supported by the Healthy Living as a Service project (with project number 13948 of the research programme KIC which is (partly) financed by the Dutch Research Council (NWO)). We would also like to thank the facilities, like the Habrok cluster provided by CIT of the University of Groningen, and other support.